OpenStudy (anonymous):

how much will the water temperature change if 4000.0j of energy is removed from a 50.0g sample of copper? Cu=0.385j/gc

OpenStudy (frostbite):

Same equation as in last question: \[\Large Q=m \times c \times \Delta T\] Solve for the change in temperature: \[\Large \Delta T=\frac{ Q }{ m \times c }\] Plug in the data and evaluate.

OpenStudy (anonymous):

thanks I got it.

OpenStudy (frostbite):

No problem.